Mandy is twice as old as Jessica. Three years ago Mandy was three times as old as Jessica was then. Find the age of each girl now
Answer by fractalier(6550)

Call their ages now m and j. Thus
m = 2j
Three years ago,
m-3 = 3(j-3)
Now substitute for m and get
2j-3 = 3(j-3)
2j-3 = 3j-9
Now subtract 2j and add nine to both sides and get
6 = j
m = 12